### v3

@@ 5 Importance-based Selection
This section provides the review of methods on importance measurement and selection. By importance we mean the necessity of adding one instruction-response sample into the training set. Due to the pre-training nature of LLMs, a wide range of materials have been “parameterized" as internal knowledge and therefore several common tasks can be correctly solved without additional fine-tuning. In this case, alignment is not required for easy samples but becomes indispensable for difficult ones. The selected datapoints provide supplementary knowledge to activate the pre-trained LLMs on following complex instructions.
@@ 5.1 Hand-crafted Indicators
The readability indices Young and Shishido (2023) can be used to assess both quality (see Sec. §3.1) and difficulty of text samples. Specifically, samples with intricate grammar, advanced vocabulary, and inference dependency are deemed as difficult ones and can be used to evaluate robustness of models across benchmarks of various difficulty levels Smith and Johnson (2020); Kiela et al. (2021); Ethayarajh et al. (2022); Belinkov and Glass (2019); Nie et al. (2019); Ribeiro et al. (2020). For specialized domains such as solving maths problems, the education level (e.g., elementary-level, high school-level, and university-level) determines the difficulty of samples Patel et al. (2021); Huang et al. (2016); Koncel-Kedziorski et al. (2016).
One of the pioneering studies on readability scores for difficulty assessment is to compute the percentage of difficult or easy words in one sentence Klare (1974); Begeny and Greene (2014). The words on a pre-defined list are counted as familiar words, and those not listed are unfamiliar, advanced words. Besides, the average number of syllables per word, the number of single-syllable words, and the number of multi-syllable words are also indicative in assessing the text materials Connatser (1999); Carrell (1987); Zakaluk and Samuels (1988); Dale and Chall (1949). Notably, there exist three representative readability metrics: 1) the Dale Chall formula Chall and Dale (1995), 2) the flesch reading ease Flesch (1948), and 3) the gunning fog index Gunning (1952). Given these metrics, Saranathan et al. conducts a thorough analysis on existing NLP datasets $S$ to select the most challenging subsets for efficient evaluation of LLMs. The easiest and hardest samples from the TruthfulQA Lin et al. (2021) via these indicators are confirmed positively correlated with the actual complexity. The selection of difficult instruction-response pairs via Eq. 5 allows the wider performance distribution of models under investigation, making it accurate to keep the relative rank of different models unchanged on subsets $S_{b}$.
@@
The computing of difficulty indices helps comprehensively analyze the robustness of models across samples and datasets. In addition, it also presents guidelines in curating and constructing discriminating NLP benchmarks.
@@ 5.2 Model-based Indicators
To avoid potential confusion, the model-based importance indicators discussed in this section are mainly categorized as three kinds: 1) uncertainty-based; 2) reward score-based; and 3) data model-based. Methods that employ training/inference losses, errors (metrics), and gradients, despite their involvement of the language model for importance sampling, are not included.
@@
Inspired from uncertainty indicators Siddhant and Lipton (2018); Kung et al. (2023); Nieth et al. (2024) proposes the prompt uncertainty, which measures the disagreement of model responses on different perturbed versions of the same instruction:
@@
Jiang et al. (2023b) targets at the over-confidence problem of LLMs after instruction tuning Kadavath et al. (2022), and proposes to calibrate the uncertainty with augmented prompt ensembles. It casts the uncertainty estimation of either discriminative or generative tasks into a multiple-choice selection problem. Specifically for open-generation tasks, different candidate responses are designed to be as diverse as possible by: 1) prompting explicitly to encourage semantically distinct answers, or 2) clustering sampled responses (with a high temperature) into groups and choosing the prototype response from each group. Such calibrated uncertainty can be used to precisely choose important samples.
Apart from the uncertainty, the reward model can also be used beyond quality scorer. Since most of the knowledge and capabilities are acquired during pre-training Zhou et al. (2024a), the instruction tuning datasets are aimed at aligning the behavior of models with human preference and expectations. Therefore, for any given instruction $x_{i}$, if the generated response is of high quality, then the necessity of fine-tuning on this instruction is low. Accordingly, $x_{i}$ is deemed as “unimportant" and will not be chosen into the subset. In that case, the language model parameterized as $\theta$ is first prompted with $x_{i(<t)}$ to generate the response $\hat{x}_{i(\geq t)}^{\theta}$. Then, a reward model parameterized as $\phi$ acts as a necessity evaluation model:
@@
Another series of model-based importance estimation methods are based on datamodels Ilyas et al. (2022); Park et al. (2023); Jain et al. (2023); Kang et al. (2024); Chhabra et al. (2024); Saunshi et al. (2022); Ye et al. (2024), where the contribution of each datapoint to the model’s behavior is estimated. The datamodels can be implemented in any machine learning model which targets at predicting the influence of each datapoint on the performance of the trained model Koh and Liang (2017); Jain et al. (2022); Liu et al. (2024b); Picard et al. (2024); Bae et al. (2024); Covert et al. (2024).
Engstrom et al. (2024) proposes to use datamodels to select subsets that maximize the overall performance. Specifically, it chooses the subset $S_{b}\subset S,S=\{x_{1},x_{2},...,x_{|S|}\}$ by estimating the loss of the model trained on it. Out of simplicity, the datamodel $\tau_{\theta_{x}}$ can be implemented as a linear model and it learns to approximate the actual loss via the TARK estimator Park et al. (2023):
@@
where $\mathcal{L}_{x_{j}}(S_{i})$ denotes the loss of the model (trained on $S_{i}$) on the sample $x_{j}$. The $\hat{\mathbbm{E}}^{(m)}$ is a $m$-sample empirical expectation and $L_{reg}(\cdot,\cdot)$ is a regression loss function (e.g., mean squared error). Intuitively, what the datamodel $\tau_{\theta_{x}}$ does is to approximate the real loss $\mathcal{L}_{x_{j}}(S_{i})$ under various compositions of subset $S_{i}\sim S_{b}$. Given any subset $S_{b}$, the averaged loss approximated by the datamodel on all $x_{j}\in S_{eval}$ is calculated on the evaluation set $S_{eval}$ and minimized to find the optimal $S_{b},|S_{b}|=b$:
$S_{b}^{*}$ $=\operatorname*{arg}\operatorname*{min}_{S_{b}\subset S}\hat{\mathcal{L}}_{S_{eval}}(S_{b}),$ (45)
^$\hat{\mathcal{L}}_{S_{eval}}(S_{b})$ $=\hat{\mathbbm{E}}^{(n)}_{x_{j}\sim S_{eval}}[\tau_{\theta_{x_{j}}}(\mathbbm{1}_{S_{b}})]$
$=\frac{1}{|S_{eval}|}\sum_{x_{j}\in S_{eval}}\theta^{T}_{x_{j}}\mathbbm{1}_{S_{b}}$
$=\mathbbm{1}_{S_{b}}^{T}(\frac{1}{|S_{eval}|}\sum_{x_{j}\in S_{eval}}\theta_{x_{j}}).$
@@
Liu et al. (2024b) also proposes a simulence-based Guu et al. (2023) linear datamodel that correlates the training samples with the validation or test set loss. A featurized simulator, namely GPTfluence, models the training dynamics (e.g., loss, BLEU and ROUGE scores) across time via an $n$-th order Markov process. It extracts representations $g(x_{i}),x_{i}\in S$ from BERT or GPT, and generates both multiplicative and additive factors to reflect the influence of any training example on the testing set. The testing performance $\phi_{t}$ at any time $t$ is affected by: 1) its performance at preceding $n$ times and 2) the current training batch $c_{t}$:
@@
where $\mathbf{W}^{T}_{(j)},\mathbf{U}^{T}_{(j)},\mathbf{W^{\prime}},\mathbf{U^{\prime}}$ are learnable weights which are optimized by minimizing $\sum_{t=1}^{T}(y_{t}-\phi_{t}(x_{k}))^{2}$ with $y_{t}$ being the ground-truth metric score monitored during training at step $t$. The $\langle\cdot,\cdot\rangle_{F}$ denotes the Frobenius inner product. Based on the datamodel, samples that reduce evaluation loss the most are selected as influential data.
Instead of performing off-line data selection, Yu et al. (2024) proposes MATES where a small datamodel continuously selects the most effective subset for the current training of the LLM. The datamodel is updated alternatively, like a partner, to adapt to the constantly changing data preferences of the model under development.
Unlike previous datamodels that predict the influence of datapoints on the testing performance of the model, Xie et al. (2023) proposes the DSIR with importance scores estimated by the distributional resemblance. It simply assumes that training samples that resemble the evaluation set are important, and these datapoints should be selected with higher probability. Given the hashed n-grams features $h(x_{i})\in\mathbbm{N}^{m}$ of $x_{i}$, its importance score $w_{i}$ is calculated as:
@@
where $S$ and $S_{eval}$ respectively denote the training set and the evaluation set. Given the budget $b$, the subset $S_{b}$ is obtained by importance-weighted sampling without replacement $b$ times. Zhang et al. (2023b) also proposes to use a independent-cascade diffusion model Li et al. (2018); Du et al. (2014) to mimic the information diffusion process upon a directed graph on embeddings of datapoints. The most influential datapoint are selected for annotation and serve as in-context learning examples for LLMs.
@@
Compared with uncertainty and reward score, datamodel-based importance indicators are more correlated with the downstream performance since the task-specific evaluation set is introduced to provide feedback for the selection scheme.
@@ 5.3 Loss and Error-based Coreset Sampling
During training, samples that contribute more to the loss or cause worse performance are considered more important. Compared with the datamodels, the influence of each datapoint is also measured in the loss and error-based coreset sampling but differs in that such measurement is performed with the same LLM under development rather than a specifically designed datamodel.
@@
The easy samples with $Unforget_{i}=1$ can be simply discarded and the important subset $S_{b}=\{x_{i}|Unforget_{i}=0,x_{i}\in S\}$ is selected for training. Recent studies on both pre-training and instruction tuning have investigated the effectiveness of using the forgetting score for efficient data pruning Sorscher et al. (2022); Paul et al. (2021); Zhang et al. (2023a); Jin and Ren (2024a); Maini et al. (2022).
In contrast to the term “forgetting", researchers introduce the concept “memorization" Feldman (2020); Tirumala et al. (2022); Antoniades et al. (2024) for analysis on the generalization of deep models Zhang et al. (2021). The memorization of training samples is necessary for reducing close-to-optimal generalization error especially when a long-tailed disttribution is observed for the training set Feldman (2020). Specifically, the amount of label memorization on the instruction-response pair ($x_{i(<t)},x_{i(\geq t)}$) is defined as follows:
@@
where $\theta^{S}$ and $\theta^{S\backslash x_{i}}$ respectively refer to the language model parameters optimized with the entire set with and without $x_{i}$. Accordingly, the influence Feldman and Zhang (2020) of the sample $x_{i}$ on other samples $x_{k},x_{k}\neq x_{i}$ can be defined as:
@@
where $x_{k(<t)}$ and $x_{k(\geq t)}$) respectively denote the instruction and response part of $x_{k}$. In practice, the memorization and influence scores are approximated via batch-wise sampling where $N$ batches $B_{1},B_{2},...,B_{N}$ are sampled from $S$ with $|B_{i}|=n$. For each batch $B_{i}$, a language model parameterized as $\theta^{B_{i}}$ is trained to compute the memorization and influence scores of each sample $x_{i}$. It is noted that some batches contain $x_{i}$ and the others do not. Therefore, the two probability terms in Eqs. 49 and 50 are respectively averaged over multiple probability outputs of the models trained on batches with and without $x_{i}$. Sorscher et al. (2022) confirms that memorization scores (Eq. 49) demonstrate stronger performance on pruning the dataset into a significantly smaller subset $S_{b}$ than random sampling, EL2N (Eq. 10), and influence scores (Eq. 50). Suzuki et al. (2023) and Schoch et al. (2023) also follow  Feldman and Zhang (2020) to select the high-quality influential subset for LLM training.
Furthermore, Chen et al. (2024b) uses the evaluation loss to check whether the current task requires certain skills or capabilities that can be obtained by learning from the prerequisite tasks. For each task, it selects the skill-dependent datapoints that reduce evaluation loss. Mishra and Sachdeva (2020) proposes a rather simple method that adopts a proxy model (e.g., logistic regression and SVM) to train on the randomly selected subset $S_{b}$ and evaluate on the remaining set $S\backslash S_{b}$. Such process iterates over multiple times to ensure that each sample is at least validated once. The probability of each sample being correctly predicted is used as importance measurement. Likewise, Lin et al. (2022) also quantifies the average marginal effect (AME) as influence of $x_{i}$. It can be viewed as a variant of shapley value Jia et al. (2019); Ghorbani and Zou (2019); Schoch et al. (2023); Kwon and Zou (2021). Different subsets are randomly sampled to train multiple submodels and each submodel is evaluated for jointly estimating the AME via LASSO regression Lecué and Mendelson (2018).

### v4

@@ 5 Importance-based Selection
This section provides the review of methods on importance measurement and selection. By importance we mean the necessity of adding one instruction-response sample into the training set. Due to the pre-training nature of LLMs, a wide range of materials have been "parameterized" as internal knowledge and therefore several common tasks can be correctly solved without additional fine-tuning. In this case, alignment is not required for easy samples but becomes indispensable for difficult ones. The selected datapoints provide supplementary knowledge to activate the pre-trained LLMs on following complex instructions. Generally, the evaluation function $q(x_{i})$ in the context of importance can be alternatively implemented as: 1) the complexity of the datapoint itself $q_{C}$ that can be perceived by the task definition and the model’s uncertainty, 2) the contribution of each datapoint to the overall performance $q_{P}$ that can be reflected in the losses and errors during learning, and 3) the degree of gradient matching $q_{G}$ that indicates influential datapoints for effective selection. Consequently, we can unify the importance-based measurements as:
$q(x_{i})=f_{i}(q_{C}(x_{i}),q_{P}(x_{i}),q_{G}(x_{i})),$ (46)
where $f_{i}$ denotes the aggregation function. It is noted that most existing methods simply choose one of the above evaluation implementations and therefore $f_{i}$ can be designed as an exclusive choice function.
@@ 5.1 Hand-crafted Indicators
The readability indices Young & Shishido (2023) can be used to assess both quality (see Sec. §3.1) and difficulty of text samples. Specifically, samples with intricate grammar, advanced vocabulary, and inference dependency are deemed as difficult ones and can be used to evaluate robustness of models across benchmarks of various difficulty levels Smith & Johnson (2020); Kiela et al. (2021); Ethayarajh et al. (2022); Belinkov & Glass (2019); Nie et al. (2019); Ribeiro et al. (2020). For specialized domains such as solving maths problems, the education level (e.g., elementary-level, high school-level, and university-level) determines the difficulty of samples Patel et al. (2021); Huang et al. (2016); Koncel-Kedziorski et al. (2016).
One of the pioneering studies on readability scores for difficulty assessment is to compute the percentage of difficult or easy words in one sentence Klare (1974); Begeny & Greene (2014). The words on a pre-defined list are counted as familiar words, and those not listed are unfamiliar, advanced words. Besides, the average number of syllables per word, the number of single-syllable words, and the number of multi-syllable words are also indicative in assessing the text materials Connatser (1999); Carrell (1987); Zakaluk & Samuels (1988); Dale & Chall (1949). Notably, there exist three representative readability metrics: 1) the Dale Chall formula Chall & Dale (1995), 2) the flesch reading ease Flesch (1948), and 3) the gunning fog index Gunning (1952). Given these metrics, Saranathan et al. conduct a thorough analysis on existing NLP datasets $S$ to select the most challenging subsets for efficient evaluation of LLMs. The easiest and hardest samples from the TruthfulQA Lin et al. (2021) via these indicators are confirmed positively correlated with the actual complexity. The selection of difficult instruction-response pairs via Eq. 8 allows a wider distribution of performance for the models under investigation. The introduction of difficult datapoints into the subset $S_{b}$ helps keep the relative rank of different models unchanged compared with that measured on the entire set $S$.
@@
The difficulty indices help comprehensively analyze the robustness of models across samples and datasets. In addition, it also presents guidelines of curating and constructing discriminating NLP benchmarks.
@@ 5.2 Model-based Indicators
To avoid potential confusion, the model-based importance indicators discussed in this section are mainly categorized as three kinds: 1) uncertainty-based; 2) reward score-based; and 3) data model-based. Methods that employ training/inference losses, errors (metrics), and gradients are not included despite their involvement of the language model for importance sampling.
@@
Inspired from uncertainty indicators, Siddhant & Lipton (2018); Kung et al. (2023); Nieth et al. (2024) propose the prompt uncertainty which measures the disagreement between model responses on different perturbed versions of the same instruction:
@@
Jiang et al. (2023b) target at the over-confidence problem of LLMs after instruction tuning Kadavath et al. (2022), and propose the CAPE to calibrate the uncertainty with augmented prompt ensembles. They first transform all discriminative and generative tasks into multiple-choice problems, and use the LLM’s predicted probabilities over answer choices (e.g., A, B, C) for uncertainty estimation. Then, prompt augmentations are performed via paraphrasing of the template, permutation of the in-context examples and the answer choices. Multiple predictions over the answer choices are collected with the augmented prompts as inputs, which helps calibrate the uncertainty in an ensemble manner. Such calibrated uncertainty tells if an instruction-tuned LLM simply memorizes the response to a given prompt rather than truly understanding the instruction. Therefore, it can be used to precisely choose important datapoints with high uncertainty.
Apart from the uncertainty, the reward model can also be used beyond quality scorer. Since most of the knowledge and capabilities are acquired during pre-training Zhou et al. (2024a), the instruction tuning datasets are aimed at aligning the behavior of models with human preference and expectations. Therefore, for any given instruction $x_{i}$, if the generated response is of high quality, then the necessity of fine-tuning on this instruction is low. Accordingly, $x_{i}$ is deemed as "unimportant" and will not be chosen into the subset. In that case, the language model parameterized as $\theta$ is first prompted with $x_{i(<t)}$ to generate the response $\hat{x}_{i(\geq t)}^{\theta}$. Then, a reward model parameterized as $\phi$ acts as a necessity evaluation model:
@@
Another series of model-based importance estimation methods are based on datamodels Ilyas et al. (2022); Park et al. (2023); Jain et al. (2023); Kang et al. (2024); Chhabra et al. (2024); Saunshi et al. (2022); Ye et al. (2024), where the contribution of each datapoint to the model’s behavior is estimated. The datamodels can be implemented in any machine learning model which targets at predicting the influence of each datapoint on the performance of the trained model Koh & Liang (2017); Jain et al. (2022); Liu et al. (2024b); Picard et al. (2024); Bae et al. (2024); Covert et al. (2024).
Engstrom et al. (2024) propose to use datamodels to select subsets that maximize the overall performance. Specifically, it chooses the subset $S_{b}\subset S,S=\{x_{1},x_{2},...,x_{|S|}\}$ by estimating the loss of the model trained on it. Out of simplicity, the datamodel $\tau_{\theta_{x}}$ can be implemented as a linear model and it learns to approximate the actual loss via the TARK estimator Park et al. (2023):
@@
where $\mathcal{L}_{x_{j}}(S_{i})$ denotes the loss of the model (trained on $S_{i}$) on the sample $x_{j}$. The $\hat{\mathbbm{E}}^{(m)}$ is a $m$-sample empirical expectation and $L_{reg}(\cdot,\cdot)$ is a regression loss function (e.g., mean squared error). Intuitively, what the datamodel $\tau_{\theta_{x}}$ does is to approximate the real loss $\mathcal{L}_{x_{j}}(S_{i})$ under various compositions of subsets $S_{i}\sim S_{b}$. Given any subset $S_{b}^{{}^{\prime}}$, the averaged loss approximated by the datamodel on all $x_{j}\in S_{eval}$ is calculated on the evaluation set $S_{eval}$ and minimized to find the optimal $S_{b},|S_{b}|=b$:
$S_{b}$ $=\operatorname*{arg}\operatorname*{min}_{S_{b}^{{}^{\prime}}\subset S}\hat{\mathcal{L}}_{S_{eval}}(S_{b}^{{}^{\prime}}),$ (50)
^$\hat{\mathcal{L}}_{S_{eval}}(S_{b}^{{}^{\prime}})$ $=\hat{\mathbbm{E}}^{(n)}_{x_{j}\sim S_{eval}}[\tau_{\theta_{x_{j}}}(\mathbbm{1}_{S_{b}^{{}^{\prime}}})]$
$=\frac{1}{|S_{eval}|}\sum_{x_{j}\in S_{eval}}\theta^{T}_{x_{j}}\mathbbm{1}_{S_{b}^{*}}$
$=\mathbbm{1}_{S_{b}^{*}}^{T}(\frac{1}{|S_{eval}|}\sum_{x_{j}\in S_{eval}}\theta_{x_{j}}).$
@@
Liu et al. (2024b) also propose a simulation-based Guu et al. (2023) linear datamodel that correlates the training samples with the validation or test set loss. A featured simulator, namely GPTfluence, models the training dynamics (e.g., loss, BLEU and ROUGE scores) across time via an $n$-th order Markov process. It extracts representations $g(x_{i}),x_{i}\in S$ from BERT or GPT, and generates both multiplicative and additive factors to reflect the influence of any training example on the testing set. The testing performance $\phi_{t}$ at any time $t$ is affected by: 1) its performance at preceding $n$ times and 2) the current training batch $c_{t}$:
@@
where $\mathbf{W}^{T}_{(j)},\mathbf{U}^{T}_{(j)},\mathbf{W^{\prime}},\mathbf{U^{\prime}}$ are learnable weights which are optimized by minimizing $\sum_{t=1}^{T}(y_{t}-\phi_{t}(x_{k}))^{2}$ with $y_{t}$ being the ground-truth metric score monitored during training at step $t$. The $\langle\cdot,\cdot\rangle_{F}$ denotes the Frobenius inner product. Samples that reduce the evaluation loss the most are selected as influential data.
Instead of performing off-line selection, Yu et al. (2024) propose MATES where a small datamodel continuously selects the most effective subset for the current training of the LLM. The datamodel, like a partner, is updated alternatively to adapt to the constantly changing preferences of the model under development.
Unlike previous datamodels that predict the influence of datapoints via the testing performance of the model, Xie et al. (2023) propose the DSIR with importance scores estimated by the distributional resemblance. It simply assumes that training samples that resemble the evaluation set are important, and these datapoints should be selected with higher probability. Given the hashed n-grams features $h(x_{i})\in\mathbbm{N}^{m}$ of $x_{i}$, its importance score $w_{i}$ is calculated as:
@@
where $S$ and $S_{eval}$ respectively denote the training set and the evaluation set. Given the budget $b$, the subset $S_{b}$ is obtained by importance-weighted sampling without replacement $b$ times. Zhang et al. (2023c) also proposes to use a independent-cascade diffusion model Li et al. (2018); Du et al. (2014) to mimic the information diffusion process upon a directed graph on embeddings of datapoints. The most influential datapoint are selected for annotation and serve as in-context learning examples for LLMs.
@@
Compared with uncertainty and reward score, datamodel-based importance indicators are more correlated with the downstream performance since the task-specific evaluation set is introduced to provide feedback on the selection scheme.
@@ 5.3 Loss and Error-based Coreset Sampling
During training, samples that contribute more to the total loss or cause worse performance are considered more important. In the light of this statement, the influence of each datapoint can also be measured via the losses and errors for coreset sampling. Compared with the datamodel-based measurement that estimates individual importance with a specifically designed datamodel, loss and error-based measurement is performed with the same LLM under development.
@@
The easy samples with $Unforget_{i}=1$ can be simply discarded and the important subset $S_{b}=\{x_{i}|Unforget_{i}=0,x_{i}\in S\}$ is selected for training. Recent studies on both pre-training and instruction tuning have investigated the effectiveness of using the forgetting score for efficient data pruning Sorscher et al. (2022); Paul et al. (2021); Zhang et al. (2023a); Jin & Ren (2024a); Maini et al. (2022).
In contrast to the term "forgetting", researchers introduce the concept "memorization" Feldman (2020); Tirumala et al. (2022); Antoniades et al. (2024) for analysis on the generalization of deep models Zhang et al. (2021). The memorization of training samples is necessary for reducing close-to-optimal generalization error especially when a long-tailed disttribution is observed for the training set Feldman (2020). Specifically, the amount of label memorization on the instruction-response pair ($x_{i(<t)},x_{i(\geq t)}$) is defined as follows:
@@
where $\theta^{S}$ and $\theta^{S\backslash x_{i}}$ respectively refer to the language model parameters optimized with the entire set with and without $x_{i}$. Accordingly, the influence of the sample $x_{i}$ on other samples $x_{k},x_{k}\neq x_{i}$ can be defined as Feldman & Zhang (2020):
@@
where $x_{k(<t)}$ and $x_{k(\geq t)}$) respectively denote the instruction and response part of $x_{k}$. In practice, the memorization and influence scores are approximated via batch-wise sampling where $N$ batches $B_{1},B_{2},...,B_{N}$ are sampled from $S$ with $|B_{i}|=n$. For each batch $B_{i}$, a language model parameterized as $\theta^{B_{i}}$ is trained to compute the memorization and influence scores of each sample $x_{i}$. It is noted that some batches contain $x_{i}$ and the others do not. Therefore, the two probability terms in Eqs. 54 and 55 are respectively averaged over multiple probability outputs of the models trained on batches with and without $x_{i}$. Sorscher et al. (2022) confirm that memorization scores (Eq. 54) demonstrate stronger performance on pruning the dataset into a significantly smaller subset $S_{b}$ than random sampling, EL2N (Eq. 13), and influence scores (Eq. 55). Suzuki et al. (2023) and Schoch et al. (2023) also follow  Feldman & Zhang (2020) to select the high-quality influential subset for LLM training.
Furthermore, Chen et al. (2024b) use the evaluation loss to check whether the current task requires certain skills or capabilities that can be obtained by learning from the prerequisite tasks. For each task, it selects the skill-dependent datapoints that reduce evaluation loss. Mishra & Sachdeva (2020) propose a rather simple method that adopts a proxy model (e.g., logistic regression and SVM) to train on the randomly selected subset $S_{b}$ and evaluate on the remaining set $S\backslash S_{b}$. Such process iterates over multiple times to ensure that each sample is at least validated once. The probability of each sample being correctly predicted is used as importance measurement. Likewise, Lin et al. (2022) also quantify the average marginal effect (AME) as influence of $x_{i}$. It can be viewed as a variant of shapley value Jia et al. (2019); Ghorbani & Zou (2019); Schoch et al. (2023); Kwon & Zou (2021). Different subsets are randomly sampled to train multiple submodels and each submodel is evaluated for jointly estimating the AME via LASSO regression Lecué & Mendelson (2018).